# A comparative analysis of metamodels for lumped cardiovascular models, and pipeline for sensitivity analysis, parameter estimation, and uncertainty quantification

### v1

@@ 1 Introduction
Many surgeries and percutaneous procedures, e.g., to treat cardiovascular diseases, organ cancer, or chronic diseases, lead to post-procedural changes in the blood circulation. These changes affect pressure and perfusion of organs and can lead to significant morbidity or mortality, but are often difficult to predict or assess based on pre-procedural metrics. Advances in the field of computational mechanics and increased computational power have made it possible to assist in personalized treatment planning or assessment by predicting said hemodynamic changes after intervention.
Three-dimensional (3D) models of blood flow dynamics and heart or vessel wall mechanics can provide high-fidelity solutions, for instance, using finite element [1] or finite volume methods [2]. Additionally, less widely used methods, like meshless approaches, including the smoothed-particle hydrodynamics [3], can also be adopted. In general, 3D simulations can be seen as a powerful tool for predicting clinically relevant biomarkers such as pressures, blood flow distribution, and wall shear stresses. However, the accuracy of 3D methods comes at the expense of high computational cost, which can reach days even on high-performance computers. This can be an issue when real-time predictions are needed, for instance, during surgery. It is also an issue for surgery optimization and sensitivity analysis problems when several simulations are required. Furthermore, these 3D models are often only feasible when modeling a specific vessel and/or organ, leaving full 3D simulations of the entire body circulation out of reach.
Reduced order models offer an attractive approach to alleviate this issue. Zero-dimensional (0D) cardiovascular models are developed to simulate the whole circulation hemodynamics [4, 5]. 0D models are based on the analogy between hydraulic and electrical systems, connecting different components of the circulation. This analogy leads to a system of differential-algebraic equations (DAE, coupled ordinary differential equations and algebraic equations), representing mass conservation, momentum balance and constitutive laws in different circulatory compartments. 0D models provide faster predictions of pressures, volumes, flow rates, and tracer concentration in the whole cardiovascular system, compared to 3D models since they only depend on solving a set of DAEs. This leads to reduced accuracy due to the simplification of the highly complex system and geometry. They can be stand-alone or provide boundary conditions for high fidelity 3D models [6].
@@
Various methods can be employed to build metamodels, with feed-forward neural networks (FFNN) being one of the most widely used approaches across different applications. FFNNs are favored for their ability to effectively approximate any continuous function when the appropriate weights (network parameters) are applied, a concept supported by the universal approximation theorem [7]. Other neural network architectures have also demonstrated significant success in specific domains; for example, convolutional neural networks (CNNs) excel in computer vision tasks [8], while recurrent neural networks (RNNs) are highly effective for time series data, particularly in natural language processing [9]. Neural networks generally perform exceptionally well when large datasets are available. However, in the context of building metamodels for hemodynamics, comprehensive and extensive patient data are often lacking. As an alternative, this database can be generated from 0D models. Despite the widespread application of neural networks in other fields, their use in developing reduced cardiovascular models is still in the exploratory stages.
@@
Despite the increased interest in the topic, few research works have been published on the application of neural networks for 0D cardiovascular models. Bonnemain et al. predicted the left ventricular (LV) systolic function from systemic and pulmonary arterial pressure signals by neural networks [10, 11]. Significant efforts have also been made to create surrogates of three-dimensional models to reduce computational costs by exploiting various techniques; these include Graph Neural Networks and Physics-Informed Neural Networks (PINNs) [12, 13, 14] or Latent Neural Networks [15].
@@
Other methods have also been employed over the years, such as polynomial chaos expansion (PCE). Originally developed to represent a random output variable in terms of a polynomial function of multiple random input variables, [16] PCE exploits the properties of orthogonal polynomials to generate a surrogate model. It is mainly popular in the context of SA, as it can be used to approximate the decomposition of variance of model output with respect to model input, thereby generating Sobolíndices [17, 18]. While PCE has been used in cardiovascular applications, it has seen more usage in sensitivity analysis [19, 20, 21] than as an emulator [22].
@@
Another type of methodology is the Gaussian Processes (GPs). GPs are a versatile non-parametric statistical framework used for regression and classification tasks, providing a systematic way to quantify uncertainty. This makes them highly useful across a wide range of fields, for instance, in the field of machine learning [23], robotics [24], bioinformatics [25], and medicine [26]. In the field of cardiovascular models, GPs have been applied to the study of ventricular mechanics [27, 28] and pulmonary circulation [29], while more recently, they have been implemented for reconstructing blood flow regimes in realistic geometries [30].
@@
In this paper, a pipeline to build surrogate models for 0D cardiovascular systems is proposed. FFNN, PCE, and GPs are assessed to build surrogates for three different cardiovascular models. The first model is a simplified whole body hemodynamics model which predicts portal vein pressure following liver surgery [31, 32]. The second one is a more complex whole body model which aids in intervention planning for Pulmonary Arterial Hypertension (PAH) patients [6]. Finally, the third model evaluates blood perfusion in an organ post-revascularization surgery via its effect on the concentration in time of an injected tracer. The first two models were selected based on their complexity, as they simulate the whole-body circulation under different clinical scenarios and generate scalar biomarkers as outputs. In contrast, the third model, though simplified, was chosen because its output of interest is a time-dependent signal, requiring a different method to build the surrogate model. More details about the three models are discussed in section 2. The application of surrogate models for conducting SA, solving inverse problems and performing UQ is also investigated as these steps are crucial to create the digital twin of a patient.
@@
Zero-dimensional models, otherwise called lumped parameter models, are reduced order models that are solely time-dependent and are typically characterized by sets of (possibly non-linear) DAEs [33, 34].
With regards to cardiovascular modelling, the key feature of the 0D models is their capability to simulate hemodynamics throughout the entire circulatory system, with a reasonable degree of accuracy. They bring the potential to offer clinicians a reliable tool for fast decision making [olse2021predicting].
@@
Table 1: Pressure-Flow relationships of the most commonly used electrical components. Note that the valve model illustrated is an ideal representation.
@@
Valve[Uncaptioned image] $p=0$ if $q>0$, $q=0$ if $p<0$
In Table 1, the non-linear resistance is a concept not often observed in 0D models, yet, in our case, it is included in the first two models. It takes into account the nonlinear effects of convective energy loss due to possible flow separation, vortices, and turbulence, mainly occurring in larger vessels [35]. Additionally, Model 1 contains a pressure-dependent resistance, which describes an increase in vessel radii due to increased pressure, thus decreasing vascular resistance [31]. Lastly, for a more dynamic representation of the valve component, the addition of a non-linear resistance and an inductance is considered, as proposed by [36], to replicate the leaflet motion. More detailed expressions for valve dynamics have also been introduced, for instance, considering regurgitation and stenosis [37, 38].
@@ 2.1 Model 1: predicting portal vein pressure after liver surgery
Model 1 aims at predicting hemodynamic changes after partial hepatectomy, whereby a significant portion of liver tissue is removed [31, 32]. Removal of tissue leads to an increase of resistance for mesenteric blood flow, which in turn will lead to an increased portal pressure. Current guidelines consider pre-operative portal hypertension (porto-caval pressure gradient $\geq$ 10 mmHg) [39] before surgery as a risk for the potential failure of the remaining liver tissue, however, [40] instead demonstrated post-operative portal hypertension (portal pressure $\geq$ 21 mmHg) to be a better predictor of post-operative liver failure. Hence, model-based prediction of portal pressure may alleviate risks of post-operative complications.
Since approximately 25% of the total cardiac output passes the liver, changes in liver resistances by removal of tissue will likely alter the total hemodynamics of the system. Therefore, the model comprises the total circulation (simplified): left and right heart chambers and associated valves, pulmonary circulation, digestive organs, two hemi-livers and remaining circulation (Figure 2). The model is personalized with pre-operative data, and resection is simulated by increasing resistance proportional to the fraction of tissue removed [31, 32].
@@
Figure 2: Schematic for Model 1, for prediction of hemodynamic change after liver resection [31, 32] It is composed of the left and right parts of the heart, pulmonary circulation, digestive organs, two hemi-livers and remaining circulation.
@@ 2.2 Model 2: pulmonary arterial hypertension intervention planning
The second model is targeted for simulating the hemodynamic conditions under Pulmonary Arterial Hypertension (PAH) [41]. PAH is a rare disease, which can be idiopathic or related to secondary diseases, and it is described by an abnormal increase in the pulmonary arterial pressure. It can lead to right ventricular overloading, progressively inducing a right ventricular failure, and if left untreated, it can be life-threatening [42]. This zero-dimensional model, originally proposed in [6], serves as a predictive tool for intervention planning. When drug therapy proves ineffective, procedures such as the creation of a pulmonary-to-systemic shunt, commonly known as Potts shunt [43], provides an alternative to lung transplantation for a specific group of patients.
In terms of its technical characteristics, the main distinction between this model and the first one outlined in this study lies in its utilization of the single fiber model [44, 45] to simulate cardiac electromechanics, as opposed to the time-varying elastance model employed in the former. In addition to that, the upper and lower body circulations are represented more finely, while the shunt region, including the ascending and descending aorta, the aortic arch, and the right and left pulmonary arteries, are replicated in detail, in contrast to the first focusing on the liver circulation. This closed-loop circulation model can also be coupled with a 3D model for performing CFD simulations in the shunt region as shown in [6], in a similar way to has already been done for several other applications within the context of cardiovascular modelling [46, 47, 48]. The 3D flow domain is also initially required to tune this 0D model as described by [49, 50].
The model can capture the hemodynamics of the hypertensive patients before and after the shunt integration between the left pulmonary artery and the descending aorta. Personalization based on preoperative patient data is a non-trivial task [6]. Several clinically relevant key parameters can then be extracted from the model to assess the condition of the patient. Some of these are pressures, systemic and pulmonary flow rates, working volumes, and ejection fractions for assessing cardiac function and post shunt integration.
@@ 2.3 Model 3: assessment of perfusion through fluorescence patterns
The third model focuses on the time variation of a contrast agent in a certain region of interest for medical imaging applications. In addition to a 0D hemodynamic model (figure 2, table 1), the transport of a substance is evaluated. The substance is injected intravenously (IV injection) and is eliminated at a constant rate (e.g. through filtration in the kidneys or metabolization in the liver). It is transported in the blood through the main circulation and reaches the organ of interest, where the fluorescence intensity is measured. A vascular anomaly compartment is inserted before the organ of interest to model an abnormal reduction of flood flow going to the organ of interest and that results in a reduction of the organ perfusion.
@@
Figure 4: Schematic diagram of the hemodynamic and transport model for assessing perfusion in the organ of interest. The top red dashed box is the key for the hemodynamics model and the bottom green box is the key to the transport model. Here the heart is modelled by a simple flow generator (see above for more precise heart models). Hemodynamics is solved before transport; parameters $\tau$ depend on the organs and the hemodynamics properties.
@@ 3.1.1 Feed-forward neural networks
where $\mathbf{z^{0}}$ and $\mathbf{z^{L}}$ are the input and output of the function, respectively. While, $\mathbf{W^{i}}$ and $\mathbf{b^{i}}$ are the parameters of each layer, known as the weights and biases, respectively. $\sigma^{i}$ is a nonlinear function use to add nonlinearity in the representation and is called the activation function. $L$ is the number of layers in the network.
@@
where $\mathbf{\hat{Y}}$ is a neural network approximation and $\mathbf{\theta}$ a vector compromising the set of function parameters (weights and biases).
@@
Recurrent neural network architectures [51] are designed to treat time-series data where the sequential nature of the data is the key in the architecture design. Moreover, some variants such as Long-Short Term Memory (LSTM) network [52] have long term dependencies and address exploding/vanishing gradients issue.
In this work, the neural networks are implemented in the Tensorflow library [53].
@@ 3.1.2 Polynomial chaos expansion
Polynomial Chaos Expansion (PCE) was originally designed by [16] to represent random variables in the form of a set of polynomials and is often utilized in global sensitivity analysis [17]. The choice of polynomial corresponds to the probability distribution of its respective model input parameter. It is possible to approximate the response $Y$ of a system $f$ to a random input $\mathbf{X}$ using the aforementioned polynomials,
@@
However, since interaction terms are likely to contribute significantly less, several truncation schemes exist [54] to reduce the number of samples required. An important novel addition is the so-called adaptive sparse PCE, which iteratively adds and removes polynomial terms based on their contribution the surrogate model’s output of interest, thereby simultaneously reducing the number of input samples required as well as reduces the chance of overfitting [55]. In this case, model quality is given by the jack-knife error, whereby the quality is determined by sequentially removing a single data point and determining the model prediction error with respect to that point.
A benefit of PCE is that the Sobols̀ensitivity indices can be derived analytically from the PCE coefficients [17], informing the user which input parameters might be fixed parameters.
In this work, we use the Python implementation of OpenTURNS, which utilises the adaptive PCE.

### v2

# A comparative analysis of metamodels for 0D cardiovascular models, and pipeline for sensitivity analysis, parameter estimation, and uncertainty quantification

@@ 1 Introduction
Many surgeries and percutaneous procedures, e.g., to treat cardiovascular diseases, organ cancer, or chronic diseases, lead to post-procedural changes in the blood circulation. These changes affect pressure and perfusion of organs and can lead to significant morbidity or mortality, but are often difficult to predict or assess based on pre-procedural metrics. Advances in the field of computational mechanics and increased computational power have made it possible to assist in personalized treatment planning or assessment by predicting previously mentioned hemodynamic changes after intervention.
Three-dimensional (3D) models of blood flow dynamics and heart or vessel wall mechanics can provide high-fidelity solutions, for instance, using finite element [fem] or finite volume methods [fvm]. Additionally, less widely used methods, like meshless approaches, including the smoothed-particle hydrodynamics [sph], can also be adopted. In general, 3D simulations can be seen as a powerful tool for predicting clinically relevant biomarkers such as pressures, blood flow distribution, and wall shear stresses. However, the accuracy of 3D methods comes at the expense of high computational cost, which can reach days even on high-performance computers. This can be an issue when real-time predictions are needed, for instance, during surgery. It is also an issue for surgery optimization and sensitivity analysis problems when several simulations are required. Furthermore, these 3D models are often only feasible when modeling a specific vessel and/or organ, leaving full 3D simulations of the entire body circulation out of reach.
Reduced order models offer an attractive approach to alleviate this issue. Zero-dimensional (0D) cardiovascular models are developed to simulate the whole circulation hemodynamics [0D-1D_review, caiazzo:hal-01955520]. 0D models are based on the analogy between hydraulic and electrical systems, connecting different components of the circulation. This analogy leads to a system of differential-algebraic equations (DAE, coupled ordinary differential equations and algebraic equations), representing mass conservation, momentum balance and constitutive laws in different circulatory compartments. 0D models provide faster predictions of pressures, volumes, flow rates, and tracer concentration in the whole cardiovascular system, compared to 3D models since they only depend on solving a set of DAEs. This leads to reduced accuracy due to the simplification of the highly complex system and geometry. They can be stand-alone or provide boundary conditions for high fidelity 3D models [pant2022multiscale].
@@
Various methods can be employed to build metamodels, with feed-forward neural networks (FFNN) being one of the most widely used approaches across different applications. FFNNs are favored for their ability to effectively approximate any continuous function when the appropriate weights (network parameters) are applied, a concept supported by the universal approximation theorem [universal_approx]. Other neural network architectures have also demonstrated significant success in specific domains; for example, convolutional neural networks (CNNs) excel in computer vision tasks [computer_vision], while recurrent neural networks (RNNs) are highly effective for time series data, particularly in natural language processing [nlp]. Neural networks generally perform exceptionally well when large datasets are available. However, in the context of building metamodels for hemodynamics, comprehensive and extensive patient data are often lacking. As an alternative, this database can be generated from 0D models. Despite the widespread application of neural networks in other fields, their use in developing reduced cardiovascular models is still in the exploratory stages.
@@
Despite the increased interest in the topic, few research works have been published on the application of neural networks for 0D cardiovascular models. Bonnemain et al. predicted the left ventricular (LV) systolic function from systemic and pulmonary arterial pressure signals by neural networks [LV_nn, LV_device_2021]. Significant efforts have also been made to create surrogates of three-dimensional models to reduce computational costs by exploiting various techniques; these include Graph Neural Networks and Physics-Informed Neural Networks (PINNs) [kissas2020machine, de2024finite, pegolotti2024learning] or Latent Neural Networks [salvador2024whole].
@@
Other methods have also been employed over the years, such as polynomial chaos expansion (PCE). Originally developed to represent a random output variable in terms of a polynomial function of multiple random input variables, [wiener1938homogeneous] PCE exploits the properties of orthogonal polynomials to generate a surrogate model. It is mainly popular in the context of SA, as it can be used to approximate the decomposition of variance of model output with respect to model input, thereby generating Sobol sensitivity indices [sudret2008global, sala2023sensitivity]. While PCE has been used in cardiovascular applications, it has seen more usage in sensitivity analysis [eck2016guide, meiburg2020uncertainty, schaferglobal] than as an emulator [heinen2020metamodeling].
@@
Another type of methodology is the Gaussian Processes (GPs). GPs are a versatile non-parametric statistical framework used for regression and classification tasks, providing a systematic way to quantify uncertainty. This makes them highly useful across a wide range of fields, for instance, in the field of machine learning [seeger2004gaussian], robotics [deisenroth2013gaussian], bioinformatics [lawrence2005probabilistic], and medicine [coveney2020sensitivity]. In the field of cardiovascular models, GPs have been applied to the study of ventricular mechanics [di2018gaussian, strocchi2023cell] and pulmonary circulation [paun2021markov], while more recently, they have been implemented for reconstructing blood flow regimes in realistic geometries [ashtiani2024reconstructing].
@@
In this paper, a pipeline to build surrogate models for 0D cardiovascular systems is proposed. FFNN, PCE, and GPs are assessed to build surrogates for three different cardiovascular models. The first model is a simplified whole body hemodynamics model which predicts portal vein pressure following liver surgery [audebert2017partial, golse2021predicting]. The second one is a more complex whole body model which aids in intervention planning for Pulmonary Arterial Hypertension (PAH) patients [pant2022multiscale]. Finally, the third model evaluates blood perfusion in an organ post-revascularization surgery via its effect on the concentration in time of an injected tracer. The first two models were selected based on their complexity, as they simulate the whole-body circulation under different clinical scenarios and generate scalar biomarkers as outputs. In contrast, the third model, though simplified, was chosen because its output of interest is a time-dependent signal, requiring a different method to build the surrogate model. More details about the three models are discussed in section 2. The application of surrogate models for conducting SA, solving inverse problems and performing UQ is also investigated as these steps are crucial to create the digital twin of a patient.
@@
Zero-dimensional models, otherwise called lumped parameter models, are reduced order models that are solely time-dependent and are typically characterized by sets of (possibly non-linear) DAEs [formaggia2010cardiovascular, shi2011review].
With regards to cardiovascular modelling, the key feature of the 0D models is their capability to simulate hemodynamics throughout the entire circulatory system, with a reasonable degree of accuracy. They bring the potential to offer clinicians a reliable tool for fast decision making [golse2021predicting].
@@
Table 1: Pressure-Flow relationships of the most commonly used electrical components. Resistance ($R$) accounts for the viscous loss effects of blood flow, while the non-linear (quadratic) resistance ($K$) captures non-linear effects caused by factors such as flow separations, stenoses, and other complex phenomena. Inertance ($L$) reflects the inertia of blood, representing the resistance to changes in flow. Compliance ($C$) describes the ability of a vessel to expand and store blood, while time-varying compliance accounts for the dynamic changes in compliance, typically replicating cardiac chamber dynamics [Haghebaert2025]. The valve allows flow ($q$) only when the pressure difference ($p$) is positive, preventing backflow. Note that the illustrated valve model is an ideal representation.
@@
Valve[Uncaptioned image]
$p=0$ if $q>0$,
$q=0$ if $p<0$
In Table 1, the non-linear resistance is a concept not often observed in 0D models, yet, in our case, it is included in the first two models. It takes into account the nonlinear effects of convective energy loss due to possible flow separation, vortices, and turbulence, mainly occurring in larger vessels [migliavacca2001modeling]. Additionally, Model 1 contains a pressure-dependent resistance, which describes an increase in vessel radii due to increased pressure, thus decreasing vascular resistance [audebert2017partial]. Lastly, for a more dynamic representation of the valve component, the addition of a non-linear resistance and an inductance is considered, as proposed by [mynard2012simple], to replicate the leaflet motion. More detailed expressions for valve dynamics have also been introduced, for instance, considering regurgitation and stenosis [pant2018lumped, pant2016data].
@@ 2.1 Model 1: predicting portal vein pressure after liver surgery
Model 1 aims at predicting hemodynamic changes after partial hepatectomy, whereby a significant portion of liver tissue is removed [audebert2017partial, golse2021predicting]. Removal of tissue leads to an increase of resistance for mesenteric blood flow, which in turn will lead to an increased portal pressure. Current guidelines consider pre-operative portal hypertension (porto-caval pressure gradient $\geq$ 10 mmHg) [liver2012easl] before surgery as a risk for the potential failure of the remaining liver tissue, however, [allard2013posthepatectomy] instead demonstrated post-operative portal hypertension (portal pressure $\geq$ 21 mmHg) to be a better predictor of post-operative liver failure. Hence, model-based prediction of portal pressure may alleviate risks of post-operative complications.
Since approximately 25% of the total cardiac output passes the liver, changes in liver resistances by removal of tissue will likely alter the total hemodynamics of the system. Therefore, the model comprises the total circulation (simplified): left and right heart chambers and associated valves, pulmonary circulation, digestive organs, two hemi-livers and remaining circulation (Figure 2). The model is personalized with pre-operative data, and resection is simulated by increasing resistance proportional to the fraction of tissue removed [audebert2017partial, golse2021predicting].
@@
Figure 2: Schematic for Model 1, for prediction of hemodynamic change after liver resection [audebert2017partial, golse2021predicting] It is composed of the left and right parts of the heart, pulmonary circulation, digestive organs, two hemi-livers and remaining circulation.
@@ 2.2 Model 2: pulmonary arterial hypertension intervention planning
The second model is targeted for simulating the hemodynamic conditions under Pulmonary Arterial Hypertension (PAH) [farber2004pulmonary]. PAH is a rare disease, which can be idiopathic or related to secondary diseases, and it is described by an abnormal increase in the pulmonary arterial pressure. It can lead to right ventricular overloading, progressively inducing a right ventricular failure, and if left untreated, it can be life-threatening [valdeolmillos2023thirty]. This zero-dimensional model, originally proposed in [pant2022multiscale], serves as a predictive tool for intervention planning. When drug therapy proves ineffective, procedures such as the creation of a pulmonary-to-systemic shunt, commonly known as Potts shunt [blanc2004potts], provides an alternative to lung transplantation for a specific group of patients.
In terms of its technical characteristics, the main distinction between this model and the first one outlined in this study lies in its utilization of the single fiber model [arts1991relation, bovendeerd2006dependence] to simulate cardiac electromechanics, as opposed to the time-varying elastance model employed in the former. In addition to that, the upper and lower body circulations are represented more finely, while the shunt region, including the ascending and descending aorta, the aortic arch, and the right and left pulmonary arteries, are replicated in detail, in contrast to the first focusing on the liver circulation. This closed-loop circulation model can also be coupled with a 3D model for performing CFD simulations in the shunt region as shown in [pant2022multiscale], in a similar way to has already been done for several other applications within the context of cardiovascular modelling [marsden2015multiscale, lagana2005multiscale, migliavacca2006multiscale]. The 3D flow domain is also initially required to tune this 0D model as described by [pant2014multiscale, pant2014methodological].
The model can capture the hemodynamics of the hypertensive patients before and after the shunt integration between the left pulmonary artery and the descending aorta. Personalization based on preoperative patient data is a non-trivial task [pant2022multiscale]. Several clinically relevant key parameters can then be extracted from the model to assess the condition of the patient. Some of these are pressures, systemic and pulmonary flow rates, working volumes, and ejection fractions for assessing cardiac function and post shunt integration.
@@ 2.3 Model 3: assessment of perfusion through fluorescence patterns
The third model focuses on the time variation of a contrast agent in a certain region of interest for medical imaging applications. In addition to a 0D hemodynamic model (figure 2, table 1), the transport of a substance is evaluated. The substance is injected intravenously (IV injection) and is eliminated at a constant rate (e.g. through filtration in the kidneys or metabolization in the liver). It is transported in the blood through the main circulation and reaches the organ of interest, where the fluorescence intensity is measured. A vascular anomaly compartment is inserted before the organ of interest to model an abnormal reduction of blood flow going to the organ of interest and that results in a reduction of the organ perfusion.
@@
Figure 4: Schematic diagram of the hemodynamic and transport model for assessing perfusion in the organ of interest. The top red dashed box is the key for the hemodynamics model and the bottom green box is the key to the transport model. Here the heart is modelled by a simple flow generator (see above for more precise heart models). Hemodynamics is solved before transport. The $h(t)$ function represents the transformation undergone by the compartment input signal through the convolution relation $c_{\mathrm{out}}=h*c_{\mathrm{in}}$. Parameters $\tau$ depend on the organs and the hemodynamics properties.
@@ 3.1.1 Feed-forward neural networks
where $\mathbf{z}^{0}$ and $\mathbf{z}^{L}$ are the input and output of the function, respectively. While, $\mathbf{W}^{i}$ and $\mathbf{b}^{i}$ are the parameters of each layer, known as the weights and biases, respectively. $\sigma^{i}$ is a nonlinear function used to add nonlinearity in the representation and is called the activation function. $L$ is the number of layers in the network.
@@
where $\mathbf{\hat{Y}}$ is a neural network approximation and $\mathbf{\theta}$ a vector comprising the set of function parameters (weights and biases).
@@
Recurrent neural network architectures [rnn] are designed to treat time-series data where the sequential nature of the data is the key in the architecture design. Moreover, some variants such as Long-Short Term Memory (LSTM) network [lstm] have long term dependencies and address exploding/vanishing gradients issue.
In this work, the neural networks are implemented in the Tensorflow library [tensorflow2015-whitepaper].
@@ 3.1.2 Polynomial chaos expansion
Polynomial Chaos Expansion (PCE) was originally designed by [wiener1938homogeneous] to represent random variables in the form of a set of polynomials and is often utilized in global sensitivity analysis [sudret2008global]. The choice of polynomial corresponds to the probability distribution of its respective model input parameter. It is possible to approximate the response $Y$ of a system $f$ to a random input $\mathbf{X}$ using the aforementioned polynomials,
@@
However, since interaction terms are likely to contribute significantly less, several truncation schemes exist [blatman2008sparse] to reduce the number of samples required. An important novel addition is the so-called adaptive sparse PCE, which iteratively adds and removes polynomial terms based on their contribution to the surrogate model’s output of interest, thereby simultaneously reducing the number of input samples required as well as reducing the chance of overfitting [blatman2011adaptive]. In this case, model quality is given by the jack-knife error, whereby the quality is determined by sequentially removing a single data point and determining the model prediction error with respect to that point.
A benefit of PCE is that the Sobol sensitivity indices can be derived analytically from the PCE coefficients [sudret2008global], informing the user which input parameters might be fixed parameters.
In this work, we use the Python implementation of OpenTURNS [OpenTURNS], which utilises the adaptive PCE.